At a particular speed, a jumbo jet can travel its own length in 20 seconds. At the same speed, the jumbo jet taxied completely past a 710 foot long hangar in 70 seconds. What is the length of the jumbo jet?
Thanks for your help.

Answer by stanbon(48558)

At a particular speed, a jumbo jet can travel its own length in 20 seconds. At the same speed, the jumbo jet taxied completely past a 710 foot long hangar in 70 seconds. What is the length of the jumbo jet?
----------------
Let p be the length of the plane in feet.
------------------
Flying DATA:
distance = p ft. ; time = 20 sec. ; rate = p/20 ft/sec
------------------
Taxiing DATA:
distance = 710+p ; time = 70 sec. ; rate = (710+p)/70
--------------------
EQUATION:
rate = rate
p/20 = (710+p)/70
70p = 14200+20p
50p = 14200
p = 284 ft
=================
Cheers,
Stan H.